Suppose 

MPC = 0.8 

t = 0.25 

Initial Equilibrium level of GDP be Y0 = 600 

∆G = 10 

∆TR = -10 

(a) Find the change in equilibrium level of GDP i.e. ∆Y0.

(b) What is the change in the Budget Surplus i.e. ∆BS? Why has BS changed?


1
Expert's answer
2021-06-15T08:59:31-0400

Solution


"\frac{∆Y}{∆G}=\frac{1}{1-MPC}"


"\frac{∆Y}{10}=\frac{1}{1-0.8} =

∆Y=50"


∆Y0=50


b.) Change in budget surplus


∆BS="\frac{(1-MP)(1-t)}{(1-MP)(1-t)} .∆G"


"∆BS=\frac{(1-0.8)(1-0.25)}{(1-0.8)(1-0.25)} .10"


∆BS= -3.75

Increase in government expenditure leads to change in equilibrium GDP thus causing a negative change in the Budget Surplus.